Two vectors of magnitudes 20cm and 10cm make angles of 120o and 150o respectively with x-axis in the xy plane. Find the magnitude and direction of their cross product.
1
Expert's answer
2018-11-05T15:46:09-0500

The angle between vectors is

θ=150-120=30°

The magnitude of their cross product:

c=ab sin⁡θ=20(10) sin⁡30=20(10)0.5=100 cm^2.

Two vectors lie in the xy-plane. Thus, their cross product oriented perpendicularly to this plane.

The direction of their cross product is z-direction.